OpenStudy (anonymous):

Without calculating determinant of matrix A, B, and C, how can I find whether [A]+[B] = [C]?

OpenStudy (mathteacher1729):

Matrix A and matrix B must have the same dimension in order for their addition to be defined.

OpenStudy (anonymous):

if they have the same dimension, what should I do next?

OpenStudy (mathteacher1729):

Is a specific dimension given?

OpenStudy (anonymous):

well u don't need to find the determinant of any matrix. u just add the corresponding elements of matrix A and matrix B and see whether u get matrix C or not. so the question of finding the determinant does not arise.

OpenStudy (anonymous):

@mathteacher1729 A=\[\left[\begin{matrix}1&2&4&0 \\1 &1&7&2\\4&2&-3&1\\5&9&2&8\end{matrix}\right]\] B=\[\left[\begin{matrix}1&2&4&0\\0&-1&3&2\\2&1&1&-1\\3&8&1&9\end{matrix}\right]\] C=\[\left[\begin{matrix}1&2&4 & 0\\0& -1&3&2\\6&3&-2&0\\5&9&2&8\end{matrix}\right]\]

OpenStudy (mathteacher1729):

This does not appear to be a determinant problem at all. Simply add the matrices and see if A + B = C ?

OpenStudy (anonymous):

nah,they're not the same but I still have to prove that [A]+[B}=[C]

OpenStudy (mathteacher1729):

I'm sorry but the problem statement is not clear. do you mean: Det(A) + Det(B) = Det(C) ? where Det(X) means The determinant of matrix X ?

OpenStudy (anonymous):

yes

OpenStudy (mathteacher1729):

You need to pay careful attention to the elementary row operations you are performing on each matrix. (Thm 1 from this link sums it up nicely) http://tutorial.math.lamar.edu/Classes/LinAlg/DeterminantByRowReduction.aspx

OpenStudy (anonymous):

okay, thank you!